# SLIM: Sim-to-Real Legged Instructive Manipulation via Long-Horizon Visuomotor Learning

### v2

@@ D-A More Dynamics Gap Reduction Techniques
Intuitively, we know that if the tracking error $\mathbf{e}_{\textrm{arm}}$ is equivalent between sim and real for all encounterable states, then there is essentially no dynamics gap. With this, one solution would be to match the tracking errors as closely as possible through accurate embodiment modeling with low-level PD control tuning. In practice, achieving this alignment for the entire state distribution proves to be nontrivial given both the nonlinearity of manipulator dynamics in conjunction with numerous unmodeled effects such as wear-and-tear, motor backlash, etc.
@@
Rather than matching non-zero tracking errors, another approach is to simply achieve near-zero tracking errors in both sim and real. In other words, whatever arm command is outputted, the resulting arm position should be roughly achieved $\mathbf{s}^{t}_{\textrm{arm}}\approx\mathbf{a}^{t-1}_{\textrm{arm}}$. In fact, prior studies have shown that tracking error minimization is one of the most reliable strategies for enabling successful sim2real transfer [3]. To do so, we use a PID joint position controller for our manipulator in both simulation and in real, with a small joint position change limit $z$ (Section III-A) to ensure minimal tracking error for a wide variety of pose configurations.
Though the leg joints also possess a dynamics mismatch, we note that policy’s task success is minimally impacted by sim-to-real locomotion differences. Therefore, we opt to simply use PD position control for the legs.
D-B More Vision Gap Reduction Techniques
@@
Vision sim-to-real, especially for colors, is notoriously difficult [41]. We simplify the problem by only using four distinct colors in our experiments, namely, red, green (more like cyan), blue and yellow. These colors are sampled from real objects in different lighting conditions and are randomized in the HSV space during simulation training. Color samples are shown in Figure A.12.
@@
Appendix E Robot Implementation Details
E-A Control Latency Management
@@
    Sensor delay randomizations: motor sensor reading delay is sampled from $(3,13)$ms per episode with $\pm$2ms variation uniformly per control step. Camera observation delay is sampled from $(30,50)$ms uniformly per step with observations taken at 60HZ. Low-level control delay is sampled from $(5,7)$ms uniformly per step. About the high-level inference delay, we evaluated the policies trained without high-level inference delay and compared the performances with and without test time high-level inference delays. We did not find clear performance differences in terms of the success rate. Therefore, we did not use high-level control delay in policy training in this work. Meanwhile, it is an option that is provided in our implementation and can be turned on whenever necessary.
Figure A.15: The architecture of the student’s policy input encoder that generates a compact observation embedding to be used by RL.
E-B Safety Measures for the Physical Robot
To reduce damage to the robot itself and the surroundings, we employ many strategies.
@@ E-B Safety Measures for the Physical Robot
    Bridging the sim-to-real gap is one important part of it, so we can have confidence after observing the robot’s behavior in simulation.
@@
    PID arm control, with reasonable action clipping, results in much smoother and more reliable arm motion, and reduces damage to the arm motors significantly.
@@
    The randomization of motor strength, payload, center of mass, arm mounting position, etc. allows the policy to adapt to the real world more easily, resulting in very accurate grasping execution, so the arm does not push into the ground or the target objects, which can lead to motor damage.
@@
    Control delay caused by policy inference time and communication delay leads to poor low level control. Go1 can topple over, and damage components mounted on top. For example, high level inference can sometimes take over 40ms on the laptop. We use asynchronous inference of high and low level to make sure high level inference does not prevent the robot to receive low level commands. We find that the latency of the Go1 Wi-Fi module can spike to 100ms at times, and decide to rely on cable connection instead. We also allow terminating deployment with a button press. When latency still spikes, e.g. when the deployment machine suddenly freezes or network cable disconnects, our driver onboard Go1 automatically stops the quadruped’s motion. This stopping is done using a scripted policy, and is not always safe, but it prevents damage to the hardware in most cases.
@@
    Additionally, to prevent the robot from wandering off and bumping into the surroundings, during training, we incorporate a rotational search behavior as explained in Section A-A during policy training, for all baseline methods and the proposed SLIM method.
Appendix F Real-World Experiment Details
F-A Containers and Graspable Objects
We use cubes of different colors as graspable objects, and baskets of different colors as the target container for dropping into, as show in Figure A.16. The size of the basket is about $16{\rm cm}\!\times\!16{\rm cm}\!\times\!16{\rm cm}$. The size of the cube is about $2.5{\rm cm}\!\times\!2.5{\rm cm}\!\times\!2.5{\rm cm}$ and the weight of the cube is about $9{\rm g}$.
@@ F-A Containers and Graspable Objects
F-B Real-world Evaluation Protocols
We detail the protocols used for real-world quantitative evaluation in the main paper (Table I).
TABLE VI: Scene Setups for Real-World Quantitative Evaluations. Legends: $\uparrow$ denotes the position robot with its orientation aligned with the direction of $\uparrow$. Uppercase letters denote baskets and lower case letters denote the objects to be grasped. The color of an object is denoted both by the background color and the character (e.g. g denotes the green cube and R denotes the red basket.)
@@ F-B Real-world Evaluation Protocols
    To ensure the repeatability of the scene setup, we use the Standard spatial layout as shown in Figure 8, as it is easy to ensure accuracy of the scene setup in terms of spatial positions. In practice, we mark the positions for the robot at the center of the square. Similarly, we mark the 4 corner positions on the lobby floor for placing the objects and baskets. The size of the square is about $2\times 2$ meters.
@@
    We follow a scene initialization procedure as detailed in Table VI across the 20 episodes to ensure good coverage of the relative positioning of the objects, the target object and basket in the language instruction, and the colors of the objects.
F-C Visualization Predicted Segmentation and Depth Maps
@@
In Figure A.17, we show the robot’s prediction results for depth and object segmentation, during an arbitrary outdoor deployment session. We emphasize again that the robot was deployed in this zero-shot scenario never modeled by our simulator, and no model finetuning was performed. Despite this, we can see that the prediction results are high-quality, given only a very short history ($0.5$ seconds) of a single wrist-mounted RGB camera and the robot proprioceptive state. While searching for the target cube, the perception model had to locate it with just a handful of pixels (first row). Right after the gripper holding and lifting the cube above the ground, the model correctly predicted that the target had switched from the cube to a basket (cube’s green color turned red; third row). Moreover, while searching for the target basket in the outdoor environment, the cluttered background did not confuse the depth and segmentation prediction (fourth row). The perception model was also able to correctly predict the segmentation mask for the robot itself (gripper fingers and quadruped in orange; fifth row). Overall, we find that in practice, this high-quality visual information bottleneck not only greatly improves the generalization ability of our RL policy, but also facilitates our debugging of the entire robotic system as it separates the errors of the perception model from those of the policy.

### v3

@@ E-A Dynamics Gap Reduction Techniques
Prior works [16] have used PD control for the arm joints. PD control needs to have large tracking errors, the difference between the target and the current joint positions, to compensate for lifting the arm and any grasped object against gravity.
@@
Intuitively, we know that if the tracking error $\mathbf{e}_{\textrm{arm}}$ is equivalent between sim and real for all visited states, then there is essentially no dynamics gap. With this, one solution would be to match the tracking errors as closely as possible through accurate embodiment modeling with low-level PD control tuning. In practice, achieving this alignment for the entire state distribution requires nontrivial work, given the nonlinearity and drifting of motor dynamics due to effects such as motor backlash, wear-and-tear, etc.
@@
Rather than matching non-zero tracking errors, another approach is to simply achieve near-zero tracking errors in both sim and real. In other words, whatever arm command is outputted, the resulting arm position should be roughly achieved $\mathbf{s}^{t}_{\textrm{arm}}\approx\mathbf{a}^{t-1}_{\textrm{arm}}$. In fact, prior studies have shown that tracking error minimization is one of the most reliable strategies for enabling successful sim-to-real transfer [3]. To do so, we use a PID joint position controller for our manipulator in both simulation and in real, with a small joint position change limit $z$ (Section III-A) to ensure minimal tracking error for a wide variety of pose configurations. Our tuned PID controller executes smoother and slower trajectories than those from the PD controller, allowing safe and accurate deployment.
We stay with PD control for the quadruped motors as the hardware driver does not support PID.
Arm Control Perturbation.
We add ±$\pm 0.02$ noise to the arm joint position targets uniformly randomly, sampled every 50 high-level control steps or 5 seconds real time. This noise is applied to both teacher and student training. The noise level is increased to ±$\pm 0.05$ after 3M iterations for teacher training. This is to make the learned policy to be robust w.r.t. different sources of control noise arising when transferred to real.
Arm Mount Perturbation.
We randomly perturb the arm mount position and yaw at the beginning of each episode to make the learned policy robust to sim-to-real gaps that may arise from the actual mounting accuracy, torso height variations, etc. In this work, we sample perturbations for $x$, $y$, $z$ and yaw $\theta$ from the following ranges: $x:\pm 1$cm, $y:\pm 1$cm, $z:\pm 1$cm, $\theta:\pm 2\times 10^{-3}$ rad.
Object Perturbations.
Due to the relatively deterministic nature of simulations (in contrast to the high-entropy real world), we observed that policies often converge to deterministic strategies. For example, for picking up an object, a policy may learn to always position itself so that the object has the same relative position to its base. This deterministic “memorizing” behavior can result in task failure if the conditions for success are even slightly off. Therefore, we enforce learning robust, reactive policies by perturbing objects of interest during various stages of task execution. Empirically, this helps with both the exploration (as the probability of stuck in a local state is reduced) and the robustness of the policy (as the state coverage is enlarged).
We perturb the object within circle with a radius of $r_{\rm perturb}$ centered around the current position of the object. We randomly sample a position within this circle and check its eligibility. If it touches or within the space of robot torso projected on to the ground plane, then this perturbation position is regarded as invalid. Otherwise it is regarded as valid. If a valid perturbation is found within 10 trails, then we transport the object to the new location as a form of object perturbation. Otherwise, we keep the position of the object unchanged. In this work, $r_{\rm perturb}\!=\!10$cm is used.
Stationary Manipulation.
We make the quadruped stand as still as possible when manipulating objects, i.e., during grasp or drop into. This avoids the large sim-to-real gap of a walking and shaking dog when accurate manipulation is needed. To achieve this, when training low-level policies to follow locomotion commands, we also encourage the quadruped to enter a stationary mode with a balanced stance when the speed command is close to zero. Without stationary manipulation, not only will success rates suffer, but also it becomes unsafe to deploy, because the gripper fingers can hit the ground, damaging fingers, arm motors and the surrounding.
Reducing Quadruped Shaking
When transitioning from locomotion to stationary, the quadruped after adding the arm on top can shake, sometimes quite violently, before settling. This only happens in real deployment, not in simulation. Our experiments seem to show that hiding the arm command and observations from the low-level locomotion policy during training and deployment reduces most of the shaking. Shaking when transitioning to stationary is still present in most low-level policies. For the experiments in this paper, we ran low level training for ten random seeds, and chose three of the best ones (less shaking and more stable stationary pose) to continue high level policy training and evaluation.
Dynamics Sim-to-Real Gaps Not Addressed.
Unlike prior work, we did not explicitly model the motor dynamics, and thus the effects of backlash and motor operating near torque limit boundaries are not modeled. For the leg motors of Go1, we follow existing work which uses P control plus damping to mimic the physical PD controller. Damping is applied on the absolute joint velocity of the motor, while the D factor in PD control is applied on the difference between the targeted and the actual joint velocities. During fast leg joint movements, this difference can cause a large sim-to-real gap.
E-B Vision Gap Reduction Techniques
The second source of sim-to-real gap for image-conditioned policies is the vision gap. This gap is caused by a visual distribution between simulated and real world pixels. Since we do not assume knowing the target scenes in advance, we have to ensure that the perception model is able to handle a wide range of vision scenarios. Accordingly, we apply various vision sim-to-real gap reduction techniques to student training, in addition to the visual information bottleneck in Section V
Image Domain Randomization.
Random Textures.
We spawn the robot on a flat ground plane and randomize the ground texture. The texture can be arbitrary and generally has no specific semantics, as long as they provide enough visual distraction to the student. This distraction will force the student to learn to focus on the most important visual features that are related to the task at hand: objects with reasonable sizes and target colors. When deployed in real, we hope that this visual focus will get transferred and help the robot locate objects.
Random Spatial Augmentation.
Random Background Objects.
To further increase the visual complexity of the surrounding environment, we randomly sample and spawn a set of background objects in the room around the robot’s initial position (Figure A.12). There are two categories of background objects: 1) primitive shapes (cuboid, cylinder, and sphere) and 2) Google scanned objects [10]. We make sure that these objects only serve the purpose of distraction and the student will not confuse them with task objects. It could collide with an background object during training with some penalty reward. When spawning an object, we also randomize its size, orientation, and texture (only for primitive shapes).
Refer to caption Refer to caption
Figure A.12: Example Training Scene with Randomized Objects. Left: isometric view; Right: birds-eye view.
@@ E-B Vision Gap Reduction Techniques
Vision sim-to-real, especially for colors, is notoriously difficult [44]. We simplify the problem by only using four distinct colors in our experiments, namely, red, green (more like cyan), blue and yellow. These colors are sampled from real objects in different lighting conditions and are randomized in the HSV space during simulation training. To obtain the randomization range, we randomly sample 1000 pixels for each color, as shown in Fig. A.13. Afterwards, we compute the means and standard deviations of the samples. HSV values are then sampled during training from a uniform distribution with a matching mean and $\pm$ bounds that cover the real-world distribution.
@@
Appendix F Robot Implementation Details
F-A Control Latency Management
@@
    Sensor delay randomizations: motor sensor reading delay is sampled from $(3,13)$ ms per episode with $\pm$ 2ms variation uniformly per control step. Camera observation delay is sampled from $(30,50)$ ms uniformly per step with observations taken at 60 HZ. Low-level control delay is sampled from $(5,7)$ ms uniformly per step.
  4. 4.
    High-level inference delay: we trained policies without high-level inference delay turned on, and compared the performance under different extra test-time high-level inference delays. We did not find clear performance differences in terms of the success rate. Therefore, we did not use high-level control delay in policy training in this work. Meanwhile, it is an option that is provided in our implementation and can be turned on whenever necessary.
F-B Safety Measures for the Physical Robot
To reduce damages to surroundings and the robot itself, we employed the following safety measures.
@@ F-B Safety Measures for the Physical Robot
    We first addressed all sim-to-real gaps of the arm alone by training and deploying a tabletop policy. Similarly, we trained and deployed a standalone quadruped policy, without the mounted arm. Only after the two standalone policies were transferred well, we started training and deploying the whole-body policy. This helped us avoid potential catastrophic failures from the first-time deployment of a complex quadruped manipulation system.
@@
    We used PID control for the arm (Section E-A), with a reasonably small action clipping, resulting in much smoother and more reliable arm movement. This strategy reduced damages to the arm motor gears significantly.
@@
    The randomizations of motor strength, payload, center of mass, arm mounting position, etc. (Section E-A), though making simulation training more difficult, actually allowed the policy to adapt to the real world more easily, resulting in very accurate grasping executions. Thus the arm did not push into the ground or the target objects due to dynamics mismatch, avoiding motor damages.
@@
    As high-level inference can sometimes take over 40ms to finish, we adopted asynchronous low-high inference (Fig. A.14) to make sure that the high-level inference did not prevent the robot from receiving closed-loop low-level commands. Otherwise, the quadruped could topple over and damage components mounted on top.
@@
    We also enabled deployment termination with a single button press. When latency spikes, especially when the deployment machine suddenly freezes or the network disconnects, our driver on Go1 will automatically stop the quadruped. This preventative stop was done using a scripted policy and it prevented damages to the hardware in most cases.
  6. 6.
    To prevent the robot from wandering off and bumping into surroundings, during training we incorporated a rotational search behavior as explained in Section A-A. This was done for all baseline methods and SLIM.
  7. 7.
    We attached a leash to the quadruped to prevent it falling down due to any reason. Later when the policy became more reliable in real, we removed the leash.
  8. 8.
    Our 3D-printed camera mount (Fig. 6) was designed to break easily, sacrificing itself to protect the camera in the event of a side fall.
  9. 9.
    During deployment, we frequently monitored the temperature of Go1’s motors and took a break if necessary. A very hot Go1 motor will result in unpredictable, dangerous locomotion behaviors.
Appendix G Real-World Experiment Details
G-A Containers and Graspable Objects
We use cubes of different colors as graspable objects, and baskets of different colors as the target container for dropping into, as show in Fig. A.15. The size of the basket is about $16{\rm cm}\!\times\!16{\rm cm}\!\times\!16{\rm cm}$. The size of the cube is about $2.5{\rm cm}\!\times\!2.5{\rm cm}\!\times\!2.5{\rm cm}$ and the weight of the cube is about $9{\rm g}$.
@@ G-A Containers and Graspable Objects
G-B Real-world Evaluation Protocols
We detail the protocols used for real-world quantitative evaluation in the main paper (Table III).
TABLE IX: Scene Setups for Real-World Quantitative Evaluations. Legends: $\uparrow$ denotes the position robot with its orientation aligned with the direction of $\uparrow$. Uppercase letters denote baskets and lower case letters denote the objects to be grasped. The color of an object is denoted both by the background color and the character (e.g., g denotes the green cube and R denotes the red basket.)
@@ G-B Real-world Evaluation Protocols
    To ensure the repeatability of the scene setup, we use the Standard spatial layout as shown in Fig. 7, as it is easy to ensure accuracy of the scene setup in terms of spatial positions. In practice, we mark the positions for the robot at the center of the square. Similarly, we mark the 4 corner positions on the lobby floor for placing the objects and baskets. The size of the square is about $2\times 2$ meters.
@@
    We follow a scene initialization procedure as detailed in Table IX across the 20 episodes to ensure good coverage of the relative positioning of the objects, the target object and basket in the language instruction, and the colors of the objects.
Appendix H Visualizations
[…]
Figure A.16: Policy Trajectory Visualization. For visualization purpose, we removed background objects. We also visualize the EE-trajectories. Trajectories corresponding to different subtasks are rendered with different colors. The ego-centric camera view is shown on the top-right of each image. The red colored arrows denote the direction of movement.
@@ Appendix H Visualizations
H-A Visualization of Policy Behaviors
The visualization of the policy behavior in real is shown in Fig. 1. We also provide demo videos in the supplementary file. Here we further visualize the policy behavior across the complete long-horizon task with the end-effector (EE) trajectory being rendered. To do this, we run the policy in the simulator, and render the images from a fixed top-down perspective and overlay the EE-trajectory. The EE-trajectory is color-coded in such a way that the segment corresponding to different subtask is rendered with a different color. The behavior of the policy is shown in Fig. A.16. To make the visualization clear we did not add background objects in the simulation for this episode. The robot start with a randomized position and yaw within a region (Start phase). A red cube and a green cube as well as a red basket and and blue basket are used. Their positions are randomized around the Standard layout and their yaw angles are also randomized. Given an instruction (“Drop the yellow cube into the red basket.”), the robot will first search for the target graspable object (the yellow cube). Note that during the Search phase, the robot actively moves both its body and the arm in searching of the target object (c.f. Fig. A.16-Search). Once the target object is in view, the robot will move towards it (Fig. A.16-MoveTo). Once the robot moves close to the target object, it will start to grasp (Fig. A.16-Grasp). During this stage, the robot body remains stationary and the arm will move towards the target cube and grasp it. After picking up the target cube, the robot will search for the target basket while keeping the cube in gripper. As shown in Fig. A.16-SearchWithObj, the robot rotates around its current position in searching of the target basket (as indicated by the circular ee-trajectory). At the same time, the arm retracts from the initial out-reaching pose after grasping towards a neutral arm pose, as indicated by the shrinking radius of the circular EE-trajectory in Fig. A.16-SearchWithObj. Once the target basket is in view, the robot will move towards it (Fig. A.16-MoveToWithObj) and drop the cube in gripper into that basket (Fig. A.16-DropInto).
H-B Visualization of Predicted Segmentation and Depth Maps
@@
In Fig. A.17, we show the robot’s prediction results for depth and object segmentation, during an arbitrary outdoor deployment session. We emphasize again that the robot was deployed in this zero-shot scenario never modeled by our simulator, and no model finetuning was performed. Despite this, we can see that the prediction results are high-quality, given only a very short history ($0.5$ seconds) of a single wrist-mounted RGB camera and the robot proprioceptive state. While searching for the target cube, the perception model had to locate it with just a handful of pixels (first row). Right after the gripper holding and lifting the cube above the ground, the model correctly predicted that the target had switched from the cube to a basket (cube’s green color turned red; third row). Moreover, while searching for the target basket in the outdoor environment, the cluttered background did not confuse the depth and segmentation prediction (fourth row). The perception model was also able to correctly predict the segmentation mask for the robot itself (gripper fingers and quadruped in orange; fifth row). Overall, we find that in practice, this high-quality visual information bottleneck not only greatly improves the generalization ability of our RL policy, but also facilitates our debugging of the entire robotic system as it separates the errors of the perception model from those of the policy.
Appendix I Human Teleoperation
We list failure modes of the Human Teleop baseline to give a more complete picture of how difficult our task is, and to show that SLIM’s performance is non-trivial.
  1. 1.
    The most common failure mode for teleoperation is task timeout. This is sometimes due to not fully grasping the cube and the cube slipping out of the fingers during movement, leaving no time for searching and grasping again. Failing to grasp fully is in turn due to either stopping a bit far from the cube, or incorrectly assessing the distance to the cube under monocular vision.
  2. 2.
    The arm can have a twisted pose due to IK near singularities. This pose either overloads the arm motors, or if not, flips the camera upside down, making the operator difficult to perceive the environment.
Overall, human teleoperation is less accurate and slower than an RL policy, due to the low-level policy’s shakiness, the difficulty of simultaneously moving the end effector during locomotion, and sometimes human errors resulting from lost focus after a long time. Besides being less time efficient, human teleoperation seems to be less energy efficient than our RL policies as well. While the RL trained high-level policy can run smoothly on a particular low-level policy, human teleoperation on the same low-level policy sometimes triggers over-current protection of Go1. This often happens when an over-extended arm is raised, causing Go1 to lose control of its leg motors. In evaluation, we ignored this Go1 hardware failure, and let the human operator simply try from start again.
Appendix J Long-Horizon Task Learning
J-A Long-Horizon Task with Bottleneck States
The challenges of long-horizon task for exploration and learning have been discussed in Section IV. Here we want to further highlight that the existence of bottleneck states, which is orthogonal to long horizon, further exacerbates the difficulties.
[…]
Figure A.18: Long-Horizon Task with Bottleneck States. Left: Top-down view of the end effector trajectories. Right: The same set of trajectories in 3D space.
A bottleneck state is a milestone that must be accomplished before reaching the final task success. Therefore, given two tasks of the same horizon length, the one with the presence of bottleneck states will typically be more difficult than the one without.
To visualizing the bottleneck states for the tasks considered in this work, we use spatial proximity of a key position (end effector/EE position) as a proxy of state proximity. Therefore, positions that are spatially close to each other can be roughly interpreted as with similar states in terms of accomplishing the mobile manipulation task under consideration. Under this setting, for the same task, if there is a common segment along the way with a compact spatial support, it can be regarded as a bottleneck state.
We visualize the trajectories from three episodes in Fig. A.18, given a fixed object layout and different starting positions. As we can see from the figure, although with quite different starting positions, all the trajectories converge and intersect at the same location marked by b in the figure, corresponding to the state for grasping the target object.
To achieve the eventual full task success (drop the target object into basket), a successful grasping of the target object has to be accomplished as a prerequisite. Given that grasping small objects requires high precision control (as shown by the inserts in the left part of Fig. A.18), this forms a clear bottleneck state, increasing the difficulties of exploration and learning, especially for the part of state space that is only available after passing the bottleneck state.
The existence of bottleneck states coupled with long task horizon pose a great challenges to standard algorithm, as shown by the results in the subsequent section.
Figure A.19: Standard Method v.s. Progressive PEX. We compare the learning behaviors of Standard method (without progressive policy expansion) and the Progressive PEX approach in terms of the task progress as training progresses.
J-B The Effectiveness of Progressive Policy Expansion
Section IV introduces the Progressive Policy Expansion approach (Progressive PEX) for long-horizon task learning. The comparison of Progressive PEX and the approach using a standard network without progressive policy expansion (Standard) on solving the long-horizon task are shown in Fig. A.19. We show each policy’s ability in solving the long-horizon task throughout the training process, measured by the number of subtasks that can be solved ($y$-axis, Long Horizon Task Progress). We order the subtasks following their respective orders appeared in the full task. We show the mean and standard deviation calculated across 3 seeds.
As can be observed from the figure, Standard approach starts with a similar trend in learning to solve subtasks as Progressive PEX. However, after learning to solve a few tasks, Standard appears to be saturated in terms of the learning progress, likely due to the inability of continual exploration and loss of plasticity [35, 36].
In contrast, Progressive PEX can keep pushing the frontier on the long-horizon task progress, learning to solve the subsequent new subtask that appears along the way towards solving the full task.